OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 - 16 = 0 and x2 = -2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

jhonyy9 (jhonyy9):

the first equation you can factoriz it x^2 -16=(x-4)(x+4) x^2 =-2x +24 x^2 +2x -24=0 factorized will be (x+6)(x-4)

jhonyy9 (jhonyy9):

can you solve these now ?

jhonyy9 (jhonyy9):

just make thses factori equal zero and will get these roots - so like x-4=0 --- x=4 and x+4=0 --- x=-4 - for the second equation can you do it the same ?

jhonyy9 (jhonyy9):

Haleybob - so you dont like coopering ?